OpenStudy (anonymous):

How would you solve this system? (Tried a bunch of times.): 5n-2m=1 4n+5m=47

hero (hero):

Solve both equations for "m"

OpenStudy (anonymous):

I got m=7.. I think it's wrong though..

hero (hero):

How did you get that?

OpenStudy (anonymous):

exactly why I need help with this one lol.

hero (hero):

Okay

hero (hero):

I'm only going to show you this once...

hero (hero):

5n-2m=1 4n+5m=47 m = (1 - 5n)/-2 m = (47 - 4n)/5 m = m 5(1-5n) =-2(47 - 4n) 5 - 25n = -94 + 8n 99 = 33n 99/33 = n n = 3 m = 7 I guess you were right :D